Problem 4-16

Future Value for Various Compounding Periods Find the amount to which $300 will grow under each of the following conditions. Do not round intermediate calculations. Round your answer to the nearest cent.

a) 11% compounded annually for 5 years

$

b) 11% compounded semiannually for 5 years

$

c) 11% compounded quarterly for 5 years

$

d) 11% compounded monthly for 5 years

$
Answer #1

Using the FV formula as:

=FV(rate,nper,pmt,pv)

a)

=FV(11%,5,,300)

=505.52

b)

=FV(11%/2,5*2,,300)

=512.44

c)

=FV(11%/4,5*4,,300)

=516.13

d)

=FV(11%/12,5*12,,300)

=518.67